OpenStudy (anonymous):

A gas occupying a volume of 725 mL at a pressure of 0.970 atm is allowed to expand at constant temperature until its pressure reaches 0.541 atm. What is its final volume?(

OpenStudy (anonymous):

I used the combined gas law for this problem and I just would like to see if I got this right.\[\frac{ P1V1 }{ N1T1 }=\frac{ P2V2 }{ N2T2 }\] I eliminated the moles and temperature since they will remain constant since we aren't changing the temperature or the gas mixture. So I am left with P1V1=P2V2 I insert the information. .970(725mL)=.541(x) I multiply together the one side and get this. 703.25=.541(x) Then I divided .541 on both sides. \[\frac{ 703.25 }{ .541 }=\frac{ .541 x }{.541 }\] which I then get 1299.9mL=x x being the V2 or second volume. Is this correct and should I convert this to Liters?

OpenStudy (matt101):

Your answer is correct and it's ok to keep mL in this case as I explain below: Generally speaking, as long as you're working in the same units on both sides, and the units scale by some factor from the "base" unit used in the equation, you're good. In the case of volume, to go from mL to L, you divide by a factor of 1000. Since you have volume on both sides, you would end up dividing both sides by 1000, so they would remain equal. Volume you do not need to convert in this case. Temperature, on the other hand, you need to be careful with. The ideal gas law is set up to use Kelvin as its unit of temperature. To go from Kelvin to Celsius, you SUBTRACT 273. To go from Kelvin to Fahrenheit, you still have to SUBTRACT 273 then multiply by 1.8 and add 32. Adding/subtracting numbers to the numerator or denominator of a fraction vastly changes the value of that fraction, and if you do it to both sides, they aren't necessarily equal. This is why you wouldn't be able to use a different unit of temperature in the equation set up as it is.

OpenStudy (anonymous):

In this case there wasn't any temperature or moles so I eliminated them from the problem all together. To just need to find the volume.

OpenStudy (matt101):

Yup you're absolutely right! I just used temperature as an example because pressure, like volume, can be scaled by various factors as well.